# GeoScaler: Geometry and Rendering-Aware Downsampling of 3D Mesh Textures

### 28 Nov 2023 v1

@@ Abstract
High-resolution texture maps are necessary for representing real-world objects accurately with 3D meshes. The large sizes of textures can bottleneck the real-time rendering of high-quality virtual 3D scenes on devices having low computational budgets and limited memory. Downsampling the texture maps directly addresses the issue, albeit at the cost of visual fidelity. Traditionally, downsampling of texture maps is performed using methods like bicubic interpolation and the Lanczos algorithm. These methods ignore the geometric layout of the mesh and its UV parametrization and also do not account for the rendering process used to obtain the final visualization that the users will experience. Towards filling these gaps, we introduce GeoScaler, which is a method of downsampling texture maps of 3D meshes while incorporating geometric cues, and by maximizing the visual fidelity of the rendered views of the textured meshes. We show that the textures generated by GeoScaler deliver significantly better quality rendered images compared to those generated by traditional downsampling methods.
@@ 1 Introduction
The two ways that texture maps differ are that they contain discontinuities and warping artifacts that arise during UV mapping, which is the process of projecting the 3D surface of a mesh onto a 2D plane. As it is often impossible to flatten a 3D surface onto a two-dimensional plane without causing major distortions, it is required to slice the 3D surfaces into a number of patches, then flatten each patch separately. Fig. 1 illustrates the trade-off between these two types of artifacts using different projections of the spherical surface of the Earth onto a 2D map. The Mercator projection (Fig. 1a) of the world map significantly distorts areas near the two poles by expanding their surface areas more than on regions closer to the equator. Conversely, the Goode homolosine projection [12] (Fig. 1b) reduces the distortions significantly but introduces significant discontinuities.
The aforementioned resampling methods operate by applying fixed-size symmetric kernels on images, to estimate the resampled values. Applying simple interpolation kernels at the discontinuities on texture maps leads to blur and/or boundary displacements. Due to the localized nature of the applied kernels, true neighbors of a pixel sitting on a discontinuity in a texture map may not be involved while filtering that pixel. Another problem arises from the symmetric nature of standard interpolation kernels, whereby pixels in all directions are weighted similarly. However, the assumption of symmetric distances is not generally true for texture maps. Textures become warped when the 3D surfaces they are painted on are flattened onto 2D planes, as the UV parametrization is always non-isometric for non-planar surfaces: areas of the triangles on the mesh in 3D are non-uniformly modified when mapped to their 2D representation on the UV plane. Ideally, the downsampling algorithm should account for warping by flexibly weighing the pixels during the interpolation
@@
Furthermore, texture maps are viewed by users in the form of rendered 2D views. Thus, it is crucial to ensure that the downsampling process is designed to optimize the perceived visual quality of the rendered content, instead of prioritizing data fidelity of the texture on the UV plane. Finally, synthesis of the UV parametrization might cause inefficient usage of space on the texture map due to the presence of a large number of discontinuities. This can lead to wasted memory consumption during rendering and may detrimentally impact the visual quality of images rendered using downsampled texture maps.
@@
    The UVWarper module computes a warping function that adjusts the UV parameters of all the vertices of the mesh to obtain a more efficient layout of the texture.
In the experiments, we demonstrate that GeoScaler significantly outperforms existing downsampling methods on a wide variety of meshes by providing up to 3dB and 2dB improvement for 8x and 4x downsampling of textures respectively, against rendered views generated from the meshes with the original textures.
@@ 2.1 Generation of Textured Meshes
When generating a texture map from textured 3D mesh surfaces, a UV parametrization process is performed which involves unwrapping the 3D model by generating a mapping from each point on its surface onto the two-dimensional UV plane. Popular UV mapping techniques include [38, 7, 3]. To reduce distortions that arise during flattening from 3D to 2D, it is common to partition each mesh into multiple patches, whereby the flattening and UV parametrization are done independently on each patch. Multiple ways of decomposing meshes into patches have been proposed [40, 26, 37].
However, noise arising during the 3D reconstruction of geometry often causes unnatural texture seams, inefficient utilization of space on the UV plane, and sub-optimal calculation of the UV mapping. These drawbacks can lead to wasted memory usage while rendering and reduced perceptual quality of the rendered images. These problems are generally exacerbated when the resolution of the texture map is reduced.
@@ 3.2 Model Architecture
The input texture map and the geometric features are then processed via an encoder-like module comprising of convolution and pooling layers which extract a set of deep features $(F_{e})$ at the desired reduced resolution of $H/s\times W/s$ and $C_{1}$ channels. This is followed by the GeoCoding module, which explicitly filters the deep features in accordance with the connectivity and layout of the mesh. Prior to reconstructing an RGB texture map, the UVWarper module estimates the warping function to be applied to the texture map and the UV parameters of the mesh vertices. The warping aims to make the texture layout and the UV mapping of the mesh more efficient. Each of these modules is explained in detail in the following sections.
@@ 3.3 GeoCoding Module
Due to the limited receptive field of convolution operations, pixels on the boundaries of a texture map are not filtered with their true neighbors on the 3D surface. The GeoCoding (short for ”geometric encoding”) module aims to address artifacts that arise due to these discontinuities of the texture map. The structure of the GeoCoding module is depicted in Fig. 3.
@@
The vertex-wise feature embeddings are filtered using residual blocks built using simple graph convolution operations as shown in Eq. 6 [35].
@@
where $N(i)$ is the set of neighbors of $i$, $e_{ij}$ is the weight between nodes $i$ and $j$, and $\theta_{1}$ and $\theta_{2}$ are learned parameters. We designate the edge weights $e_{ij}$ to equal the Euclidean distance between the mesh vertices. Performing the convolutions in the geometric domain ensures that features that are proximal in the 3D domain on the surface of the mesh, but lying on distant seams of texture maps, are appropriately fused together. The sequence of graph convolutions outputs a set of vertex-wise feature vectors of length $C_{2}$.
To reconstruct a two-dimensional map from geometrically filtered features, we first smoothly interpolate the features sampled at each vertex across the faces using barycentric coordinates. Next, the UV parametrization is again used to project features on the mesh surface onto the 2D UV plane. The output of the GeoCoding module, a two-dimensional feature map $(F_{g})$ of size $C_{2}\times H/s\times W/s$ is generated. The obtained feature set $F_{g}$ is compensated for geometric discontinuities and distortions present in the texture map. If vertices are extremely sparse, and the sampled vectors at the vertices fail to retain sufficient information, $F_{g}$ is added with a skip residual from $F_{e}$.
@@ 3.5 Rendering Loss and Optimization
$\begin{gathered}\text{RenderLoss}=\sum_{j=1}^{K}[\alpha_{1}\mathcal{L}_{L2}(I_{j}^{orig},I_{j}^{ds})\\ +\alpha_{2}\mathcal{L}_{SSIM}(I_{j}^{orig},I_{j}^{ds})+\alpha_{3}\mathcal{L}_{perceptual}(I_{j}^{orig},I_{j}^{ds})]\end{gathered}$ (8)
Given the availability of a plethora of differentiable rendering algorithms, one can be chosen that closely matches the target application. The performance of the model is validated after every few iterations by averaging the losses computed from images rendered from a set of $Q$ of poses sampled uniformly from $\mathbf{P_{K}}$.
@@ 4 Experiments
Figure 5: Qualitative results comparing GeoScaler with bicubic downsampling for 4x and 8x downsampling. The top row of meshes is from Google’s Scanned Object Dataset, the middle row is from our 3DSet5 mesh dataset, and the bottom row is from the Textured Mesh Quality Assessment dataset. For each mesh result, the top row is the Ground Truth texture, the middle row shows results from GeoScaler, and the bottom row shows results using Bicubic.
Scale
(Texture Size)
4x
8x
Dataset Mesh / Method
@@
PSNR(dB) / SSIM
Bicubic
PSNR(dB) / SSIM
@@
PSNR(dB) / SSIM
Bicubic
PSNR(dB) / SSIM
3DSet5 avocado 35.04 / 0.9559 32.48 / 0.9336 33.13 / 0.9309 29.12 / 0.8868
banjoman 41.92 / 0.9855 40.81 / 0.9823 35.18 / 0.9481 33.77 / 0.9389
bevo 30.06 / 0.9300 27.94 / 0.8982 28.68 / 0.8990 25.87 / 0.8465
clothdolls 35.96 / 0.9650 33.53 / 0.9440 32.46 / 0.9269 28.22 / 0.8580
stew 32.28 / 0.9457 29.58 / 0.9119 29.78 / 0.9041 26.17 / 0.8431
GSO shoe 28.12 / 0.9294 26.33 / 0.9155 26.07 / 0.9131 23.94 / 0.8760
shoe2 31.63 / 0.9601 30.11 / 0.9508 26.74 / 0.9118 25.28 / 0.9041
spiderman 36.53 / 0.9820 35.35 / 0.9785 33.44 / 0.9683 31.79 / 0.9606
TMQA mask 32.70 / 0.9516 31.50 / 0.9394 26.85 / 0.8631 25.46 / 0.8393
pizza 33.57 / 0.9122 32.58 / 0.8874 30.13 / 0.8306 28.54 / 0.8023
Table 1: Quantitative results comparing GeoScaler with bicubic downsampling for 4x and 8x downsampling. Each cell contains the PSNR(dB) and SSIM value. GeoScaler outperformed bicubic downsampling by a significant margin across all datasets.
Scale
GeoScaler
w/o UVWarping
w/o UVWarping &
w/o GeoCoding
4x
35.05 /
0.9564
34.45 /
0.9515
34.18 /
0.9486
8x
31.85 /
0.9218
30.54
0.8983
30.15 /
0.8895
Table 2: Ablation results of the main modules in the GeoScaler architecture. Average PSNR(dB)/SSIM obtained on the 3DSet5 dataset before and after ablating the GeoCoding and UVWarp modules.
Figure 6: Trade-off between rendering memory consumption and visual quality. Textures produced by GeoScalar produce superior visual quality while consuming the same rendering memory compared to textures produced by bicubic downsampling.
4.1 Implementation Details
Our DNN model contains 4 residual blocks in the downsampling encoder and 2 graph residual blocks in the GeoCoding module. We chose $C_{1}=C_{2}=64$, bringing the total number of parameters in the network to 0.57M. To compute the rendering loss, we used the differentiable renderer proposed in SoftRasterizer [27], along with a Soft Phong Shader [4] whose implementations are available in PyTorch3D [16]. At each iteration, we rendered $K=4$ views when computing the loss. For the rendering step, we uniformly illuminated the scene with a white ambient light. The AdamW optimizer [32] was employed for the training process. We used a peak learning rate of $35e-$ 3 with a cosine annealing schedule [31] and a warm-up over 1000 iterations. We found that it is sufficient to train the DNN over 15000 iterations for convergence. When validating the model, we used $Q=60$ poses sampled uniformly over the range of camera poses.
4.2 Datasets
We tested GeoScaler on a wide variety of textured meshes selected from multiple datasets including Google’s Scanned Object (GSO) dataset [8] and the Textured Mesh Quality Assessment (TMQA) dataset [36].
We also studied the performance of Geoscaler on unprocessed and unrefined textured meshes which are directly generated from multiple images of real-world 3D objects, which contain geometric artifacts and irregularities. To that end, we share a small dataset of five textured meshes that did not undergo any post-processing or refinement and present the performance of GeoScaler on it. We dub this dataset 3DSet5 and its meshes are shown in Fig. 4. The meshes in the 3DSet5 were generated from multiple input views (approx. 50) using Agisoft’s Metashape application. The meshes in 3DSet5 contain approximately 10000 vertices and 20000 faces and have texture maps of size 2048x2048.
4.3 Results and Ablation Study
@@
We measured the quality gains provided by GeoScaler in terms of peak signal-to-noise ratio (PSNR) and the Structural Similarity Index Metric (SSIM) as shown in Tab. 1. GeoScaler was able to significantly outperform existing methods by improving downsampling quality at both 4x and 8x scale factors. In some cases, textures obtained when performing 8x downsampling using GeoScaler produce better results than even textures obtained by 4x bicubic downsampling. This is evident on meshes avocado, bevo, and stew from the 3DSet5 dataset, as shown in Fig. 5. the objects shoe and spiderman in Fig. 5 are from Google’s Scanned Objects dataset [8], mask and pizza are from the Textured Mesh Quality Assessment dataset [36], and bevo and stew are from 3DSet5. It can be noticed that GeoScaler displays significantly superior retention of details and fine structures contained in the original texture. GeoScaler also prevents artifacts at seams on textures, as shown in 8x downsampling results of bevo (horizontal artifact in the left column of 8x bicubic result).
@@
Next, we ablated the UVWarping module and the GeoCoding module to validate their importance. Tab. 2 shows the average PSNR and SSIM values over the 3DSet5 dataset when the modules are ablated. Detailed visual results are available in the supplementary materials. In supplementary material, we also show that textures generated by GeoScaler provide perceptually better results than textures generated natively at lower resolution using MetaShape over the 3DSet5 dataset.

### 20 Mar 2024 v2

@@ Abstract
High-resolution texture maps are necessary for representing real-world objects accurately with 3D meshes. The large sizes of textures can bottleneck the real-time rendering of high-quality virtual 3D scenes on devices having low computational budgets and limited memory. Downsampling the texture maps directly addresses the issue, albeit at the cost of visual fidelity. Traditionally, downsampling of texture maps is performed using methods like bicubic interpolation and the Lanczos algorithm. These methods ignore the geometric layout of the mesh and its UV parametrization and also do not account for the rendering process used to obtain the final visualization that the users will experience. Towards filling these gaps, we introduce GeoScaler, which is a method of downsampling texture maps of 3D meshes while incorporating geometric cues, and maximizing the visual fidelity of the rendered views of the textured meshes. We show that the textures generated by GeoScaler deliver significantly better quality rendered images compared to those generated by traditional downsampling methods.
Figure 1: Qualitative comparisons of downsampling. GeoScalar can downsample texture maps of 3D meshes while providing significantly improved quality of rendered images as compared to existing methods such as bicubic resampling. The results shown here were computed on the clothdolls mesh from our 3DSet5 mesh dataset.
@@ 1 Introduction
Unlike regular 2D images, texture maps contain discontinuities and warping artifacts that arise during UV mapping, which is the process of projecting the 3D surface of a mesh onto a 2D plane. As it is often impossible to flatten a 3D surface onto a two-dimensional plane without causing major distortions, it is required to slice the 3D surfaces into several patches, and then flatten each patch separately. Fig. 2 illustrates the trade-off between these two types of artifacts using different projections of the spherical surface of the Earth onto a 2D map. The Mercator projection (Fig. 2a) of the world map significantly distorts areas near the two poles by expanding their surface areas more than on regions closer to the equator. Conversely, the Goode homolosine projection [14] (Fig. 2b) reduces the distortions but introduces significant discontinuities.
The aforementioned resampling methods operate by applying fixed-size symmetric kernels on images to estimate the resampled values. Applying simple interpolation kernels at the discontinuities on texture maps leads to blur and/or boundary displacements. Due to the localized nature of the applied kernels, true neighbors of a pixel sitting on a discontinuity in a texture map may not be involved while filtering that pixel. Another problem arises from the symmetric nature of standard interpolation kernels, whereby pixels in all directions are weighted similarly. However, the assumption of symmetric distances is not generally true for texture maps. Textures become warped when the 3D surfaces they are painted on are flattened onto 2D planes, as the UV parametrization is always non-isometric for non-planar surfaces: areas of the triangles on the mesh in 3D are non-uniformly modified when mapped to their 2D representation on the UV plane. Ideally, the downsampling algorithm should account for warping by flexibly weighing the pixels during the interpolation
@@
Furthermore, the synthesis of the UV parametrization does not take into account the content of the texture itself. This can lead to small areas on the 2D texture maps being assigned to regions with intricate color/attribute texture detail. The quality of the rendered images could be better if large but less-detailed triangles were assigned smaller areas and small but highly-detailed triangles were assigned larger areas.
Finally, texture maps are viewed by users in the form of rendered 2D views. Thus, it is crucial to ensure that the downsampling process is designed to optimize the perceived visual quality of the rendered content, instead of prioritizing data fidelity of the texture on the UV plane.
@@
    The UVWarper module computes a warping function that adjusts the UV parameters of all the vertices of the mesh to obtain a more efficient layout of the texture based jointly on a mesh’s geometry and RGB texture.
In the experiments, we demonstrate that GeoScaler significantly outperforms existing downsampling methods on a wide variety of meshes by providing up to 3dB and 2dB improvement for 8x and 4x downsampling of textures respectively.
@@ 2.1 Generation of Textured Meshes
When generating a texture map from textured 3D mesh surfaces, a UV parametrization process is performed which involves unwrapping the 3D model by generating a mapping from each point on its surface onto the two-dimensional UV plane. Popular UV mapping techniques include [49, 9, 5, 28]. To reduce distortions that arise during flattening from 3D to 2D, it is common to partition each mesh into multiple patches, whereby the flattening and UV parametrization are done independently on each patch. Multiple ways of decomposing meshes into patches have been proposed [51, 31, 48].
While these methods minimize seams and geometric distortion within texture maps, they do not pay heed to the RGB content of textures. Additionally, noise arising during the 3D reconstruction of geometry often causes unnatural texture seams, inefficient utilization of space on the UV plane, and sub-optimal calculation of the UV mapping. These drawbacks can lead to inefficient usage of space on texture maps and thus also memory usage while rendering and reduced perceptual quality of the rendered images. These problems are generally exacerbated when the resolution of the texture map is reduced.
2.2 Efficient and Alternate Texture Representation
While there has not been much focus on the downsampling of 2D texture maps, there has been significant research on compressing texture maps, rendering from compressed textures, and deriving alternate efficient representations of surface color/attributes. Space-filling 3D textures (or solid textures) were proposed in [46, 47] to counteract the detrimental effects of texture seams and warping. To combat high memory usage while rendering, the authors of [4, 24, 43, 3] proposed algorithms to build compressed representations of texture maps and directly render from them. The authors of both [41] and [44] both aim to modify standard image compression algorithms by introducing geometry-aware methods for intracoding. All of these methods can be applied along with our method to gain the benefits of both efficient compression and downsampling.
Mip ("multum in parvo") Mapping is commonly applied on texture maps to reduce aliasing when rendering meshes from a distance. This process also involves sub-sampling textures at different scales, using methods like bilinear or trilinear filtering. While some effort has been made [10, 17] to optimize mipmap generation, these techniques do not facilitate the generation high quality renders with downsampled textures at the primary viewing distance.
@@ 3.2 Model Architecture
The input texture map and the geometric features are then processed via an encoder-like module comprising of convolution and pooling layers which extract a set of deep features $(F_{e})$ at the desired reduced resolution of $H/s\times W/s$ and $C_{1}$ channels. This is followed by the GeoCoding module, which explicitly filters the deep features per the connectivity and layout of the mesh. Before reconstructing an RGB texture map, the UVWarper module estimates the warping function to be applied to the texture map and the UV parameters of the mesh vertices. The warping aims to make the texture layout and the UV mapping of the mesh more efficient. Each of these modules is explained in detail in the following sections.
@@ 3.3 GeoCoding Module
Due to the limited receptive field of convolution operations, pixels on the boundaries of a texture map are not filtered with their true neighbors on the 3D surface. The GeoCoding (short for "geometric encoding") module aims to address artifacts that arise due to these discontinuities of the texture map. The structure of the GeoCoding module is depicted in Fig. 4.
@@
The vertex-wise feature embeddings are filtered using residual blocks built using simple graph convolution operations as shown in Eq. 6 which is borrowed from [40].
@@
where $N(i)$ is the set of neighbors of $i$, $e_{ij}$ is the weight between nodes $i$ and $j$, and $\theta_{1}$ and $\theta_{2}$ are learned parameters. We designate the edge weights $e_{ij}$ to equal the Euclidean distance between the mesh vertices. Performing the convolutions in the geometric domain ensures that features that are proximal in the 3D domain on the surface of the mesh, but lying on distant seams of texture maps, are appropriately fused. The sequence of graph convolutions outputs a set of vertex-wise feature vectors of length $C_{2}$.
To reconstruct a two-dimensional map from geometrically filtered features, we first smoothly interpolate the features sampled at each vertex across the faces using barycentric coordinates. Next, the UV parametrization is again used to project features on the mesh surface onto the 2D UV plane. The output of the GeoCoding module, a two-dimensional feature map $(F_{g})$ of size $C_{2}\times H/s\times W/s$ is generated. The obtained feature set $F_{g}$ is compensated for geometric discontinuities and distortions present in the texture map. If vertices are extremely sparse, and the sampled vectors at the vertices fail to retain sufficient information, $F_{g}$ is added with a skip residual from $F_{e}$. The need for this skip residual is demonstrated more thoroughly in supplemental section Sec. 6.1
@@ 3.5 Rendering Loss and Optimization
$\begin{gathered}\text{Loss}=\sum_{j=1}^{K}[\alpha_{1}\mathcal{L}_{L2}(I_{j}^{orig},I_{j}^{ds})+\alpha_{2}\mathcal{L}_{SSIM}(I_{j}^{orig},I_{j}^{ds})+\alpha_{3}\mathcal{L}_{percep.}(I_{j}^{orig},I_{j}^{ds})]\end{gathered}$ (8)
Given the availability of a plethora of differentiable rendering algorithms, one can be chosen that closely matches the target application.
The performance of the model is validated after every few iterations by following these steps:
  1. 1.
    A set of $Q$ of poses $\mathbf{P_{Q}}$ is sampled uniformly from $\mathbb{S}$ and is predefined at the start of training.
  2. 2.
    $Q$ images ($\mathbf{I_{Q}^{orig}}$) are rendered from camera poses in $\mathbf{P_{Q}}$ using the mesh with the original high-resolution texture.
  3. 3.
    $Q$ images ($\mathbf{I_{Q}^{ds}}$) are rendered from camera poses in $\mathbf{P_{Q}}$ using the mesh with the downsampled texture.
  4. 4.
    Metrics are calculated for each pair of images in $\mathbf{I_{Q}^{orig}}$ and $\mathbf{I_{Q}^{ds}}$ and averaged over the number of elements in the set
@@ 4 Experiments
4.1 Implementation Details
Our DNN model contains four residual blocks in the downsampling encoder and two graph residual blocks in the GeoCoding module. We chose $C_{1}=C_{2}=64$, bringing the total number of parameters in the network to 0.57M. To compute the rendering loss, we used the differentiable renderer proposed in SoftRasterizer [32], along with a Soft Phong Shader [6] whose implementations are available in PyTorch3D [19]. At each iteration, we rendered $K=4$ views when computing the loss. For the rendering step, we uniformly illuminated the scene with a white ambient light. The AdamW optimizer [37] was employed for the training process. We used a peak learning rate of $35e-$ 3 with a cosine annealing schedule [36] and a warm-up over 1000 iterations. We found that it is sufficient to train the DNN over 15000 iterations to achieve convergence. When validating the model, we used $Q=60$ poses sampled uniformly over the range of camera poses.
4.2 Datasets
@@
We tested GeoScaler on a wide variety of textured meshes. We studied its efficacy on the entire Textured Mesh Quality Assessment Dataset (TMQA) [42] which consists of 55 meshes with diverse and varying geometry and color. This dataset contains meshes downloaded from SketchFab, and as such contains textured meshes generated using unknown yet varied 3D reconstruction algorithms and post-processing techniques. The curators of the dataset also portray its diversity in terms of color and geometric detail in [42]. To facilitate experimentation, we decimate the meshes to contain a maximum of 20,000 faces and texture maps of 2048x2048 resolution.
We also studied the performance of GeoScaler on unprocessed and unrefined textured meshes which are directly generated from multiple images of real-world 3D objects, which contain geometric artifacts and irregularities. To that end, we share a small dataset of five textured meshes that did not undergo any post-processing or refinement and present the performance of GeoScaler on it. We dub this dataset 3DSet5 and its meshes are shown in Fig. 5. The meshes in the 3DSet5 were generated from multiple input views (approx. 50) using Agisoft’s Metashape application. The meshes in 3DSet5 contain approximately 10000 vertices and 20000 faces and have texture maps of size 2048x2048. In the supplementary section, we analyze the performance of GeoScaler on the 3DSet5 dataset after the UV parametrizations of meshes were optimized with OptCuts[28], which reduces the number of seams and warping artifacts in the texture maps.
These testing datasets were chosen such that the proposed method is evaluated on meshes of real-world objects that were generated using a diverse range of algorithms and that underwent varied post-processing techniques.
4.3 Results and Ablation Study
Figure 6: Visual results ablating and comparing GeoScaler with existing methods for 4x downsampling. The top row of each mesh shows the results using bicubic downsampling, Lanczos downsampling, "Perceptually based downsampling" and "GS Base" from left to right. The bottom row for each mesh shows the results using "GS Base+GCM", "GS Base+UVW", "GeoScaler Full" and the original from left to right.
Figure 7: Visual results ablating and comparing GeoScaler with existing methods for 8x downsampling. The top row of each mesh shows the results using bicubic downsampling, Lanczos downsampling, "Perceptually based downsampling" and "GS Base" from left to right. The bottom row for each mesh shows the results using "GS Base+GCM", "GS Base+UVW", "GeoScaler Full" and the original from left to right.
Dataset Bicubic Lanczos PBD GS Base
GS Base
+GCM
GS Base
+UVW
@@ 4.3 Results and Ablation Study
Full
TMQA PSNR (dB) 33.72 34.11 34.24 34.97 35.42 35.55 35.88
SSIM 0.9264 0.9294 0.9286 0.9327 0.9428 0.9461 0.9524
3DSet5 PSNR (dB) 32.82 33.15 33.23 34.18 35.42 34.45 35.05
SSIM 0.9350 0.9388 0.9421 0.9486 0.9428 0.9515 0.9564
Table 1: Quantitative results of 4x downscaling comparing GeoScaler and its ablations with existing methods. GeoScaler outperforms traditional methods by a significant margin when applied to TMQA and 3DSet5 datasets.
Dataset Bicubic Lanczos PBD GS Base
GS Base
+GCM
GS Base
+UVW
@@
Full
TMQA PSNR (dB) 30.48 30.88 31.02 31.74 32.41 32.42 33.18
SSIM 0.8866 0.8935 0.8971 0.9066 0.9106 0.9115 0.9191
3DSet5 PSNR (dB) 28.63 29.34 29.53 30.15 32.41 30.54 31.85
SSIM 0.8747 0.8779 0.8812 0.8996 0.9036 0.9047 0.9218
Table 2: Quantitative results of 8x downscaling comparing GeoScaler and its ablations with existing methods. GeoScaler outperforms traditional methods by a significant margin when applied to TMQA and 3DSet5 datasets.
We measured the quality gains provided by GeoScaler in terms of peak signal-to-noise ratio (PSNR) and the Structural Similarity Index Metric (SSIM) as shown in Tab. 1 Tab. 2. GeoScaler was able to significantly outperform existing methods including bicubic downsampling, Lanczos downsampling, and "Perceptually Based Downsampling" (PBD) proposed by [45] by improving downsampling quality at both 4x and 8x scale factors. These tables also demonstrate the performance of GeoScaler after the proposed modules were ablated. The GeoScaler Base (GS Base) model is obtained by ablating both the GeoCoding module and the UVWarper Module. Its performance verifies the impact of using rendering loss and performing optimization based on the rendered space rather than the texture’s UV plane. We add the GeoCoding module to GS Base to obtain the "GS Base + GCM" model whose performance gain over GS Base validates the GeoCoding module. Similarly, "GS Base + UVW" obtained by adding only UVWarping to GS Base tests this module independently.
@@
The improvement in downsampling quality is also evident for scale factors of 4x and 8x in the visual results shown in Fig. 6 and Fig. 7 respectively. The meshes mask and pizza shown in the figure are from the Textured Mesh Quality Assessment dataset [42], while bevo and stew are from 3DSet5. It can be noticed that GeoScaler displays significantly superior retention of details and fine structures contained in the original texture. GeoScaler also prevents artifacts at seams on textures, as shown in 8x downsampling results of bevo (horizontal seam artifact in the left columns are reduced when Geocoding is applied).
Figure 8: Optimization of UV mapping by UVWarper. Highly detailed regions are assigned larger areas in texture maps produced by GeoScaler. Regions of interest marked by red boxes demonstrate the improved space allocation.
Figure 9: Trade-off between rendering memory consumption and visual quality. Textures produced by GeoScalar produce superior visual quality while consuming same rendering memory compared to textures produced using existing methods
@@
The effect of the UVWarper module is also visualized in Fig. 8 where the texture map itself is displayed. The figure shows the reassignment of areas in the texture map based on the salience of RGB details in the texture map. In the supplementary material, we also show that textures generated by GeoScaler provide perceptually better results than textures generated natively at lower resolution using MetaShape over the 3DSet5 dataset.